OpenStudy (anonymous):

suppose you roll two dice. find the number of elements in the event space of rolling a sum of 2

OpenStudy (anonymous):

can someone help me please.

OpenStudy (ganpat):

Event for sum of 2 in two dice = {2, 4 , 6, 8 , 10, 12} = 6 elements..

OpenStudy (anonymous):

how would i right tha in a fraction please.

OpenStudy (ganpat):

so even occurring = 6/12 = 1/2 where, 12.. total number on two dice (6 + 6) and 6, is no of element..

OpenStudy (anonymous):

thanks

OpenStudy (ganpat):

yup.. understood ?

OpenStudy (anonymous):

yes very